# CoDiff: Conditional Diffusion Model for Collaborative 3D Object Detection

### v2

@@ IV Our Proposed Method
CoDiff is designed as a generative model specifically multi-agent 3D object detection which is capable of simultaneously addressing pose errors and time delays. Figure 2 shows the overall architecture of CoDiff. It consists of two key ideas. Firstly, we propose a perceptual compression model to compress high-dimensional feature maps into a low-dimensional, latent space. This compression focuses on the essential semantic information of the data. Secondly, we design a conditional diffusion model that treats feature maps from different agents as conditions, combining the model’s generative capabilities with these conditions to produce more accurate and fine-grained multi-agent fused features. These features are ultimately fed into the detection head to achieve precise 3D object detection results.
IV-A Perceptual Compression Model
Most feature fusion methods employ autoregressive, attention-based transformer models, which are trained in high-dimensional feature spaces [22, 16]. However, we observe that diffusion models incur costly function evaluations during high-resolution feature sampling, presenting significant challenges in terms of computation time and energy resources. Therefore, we aim to perform sampling in a lower-dimensional space, where diffusion models can learn within this perceptually equivalent, lower-dimensional space, thereby reducing computational complexity. This approach has demonstrated its feasibility in the domain of high-resolution image synthesis [35]. Based on the above analysis, we propose a perceptual compression model that abstracts high-dimensional feature details into a lower-dimensional latent space. This space is more suitable for likelihood-based generative models because it makes computations more efficient and enabling a focus on the important semantic aspects of the data.
@@ IV-A Perceptual Compression Model
Our perceptual compression model is based on a neural network autoencoder architecture. This autoencoder comprises an encoder and a decoder, focusing on compressing and reconstructing input tensors along the channel dimension. The encoder employs a series of convolutional layers, batch normalization, and ReLU activation functions to compress high-dimensional input tensors into a lower-dimensional latent space. The decoder, using a similar structure, aims to reconstruct a low-dimensional latent representations back to the original dimensionality. This autoencoder effectively reduces computational overhead while preserving key features of the data.
@@
Precisely, given a feature map $F_{i}\in\mathbb{R}^{H\times W\times C}$, where $H,W,C$ represents height, width and channels, respectively. The encoder ${\mathcal{E}}$ encodes $F_{i}$ into a latent representation $z={\mathcal{E}}\left(F_{i}\right)$ using a compression rate $\tau$, and the decoder $\mathcal{D}$ reconstructs the feature from the latent space using the same compression rate. Therefor, the process of compressing features $F_{i}$ into the latent space is:
@@ IV-B Latent Diffusion Model
With the perceptual compression model trained in the previous step, the diffusion model can now be trained in this lower-dimensional, computationally more efficient latent space. Our goal is to generate noise-free features $F^{\prime}_{i}$ (Step (1b)) through this latent space and the decoder $\mathcal{D}$, which should represent the aggregation of features from multiple agents. We can decode the latent representation of features $F^{\prime}_{i}$ from the latent space back to the original feature space with a single pass through $\mathcal{D}$ during training.
@@
Inspired by literature [35], diffusion models can leverage conditional mechanisms $p\left(x|y\right)$ to control the target generation process, thereby improving the model’s accuracy. We find that in multi-agent collaborative perception, the information from individual agents can be naturally used as conditional inputs to generate the final fused features. Thus, the generation of multi-agent features can be controlled by manipulating the input conditions $y$.
@@
The Design of Condition. In practice, we select the feature information from $N-1$ individual agents as the condition $y_{cond}$. These informations are obtained by generating features $F_{j}$ through point cloud encoders, which are then transmitted to the Ego agent $i$, denoted as $\mathcal{M}_{j\rightarrow i}$. We strictly follow the standard DPM model to add noise, while the difference is that we employ condition-modulated denoising, which is shown in Figure 2. By progressively denoising the samples, we hope that the conditional diffusion model can act as a more flexible condition generator, aiding in learning fine-grained object information, such as precise boundaries and highly detailed shapes, especially under noisy conditions.
@@
Given noisy feature $z_{t}$ and condition $y_{cond}=\{\mathcal{M}_{j\rightarrow i}\}$ at time step $t$, $z_{t}$ is further encoded and interacts with $y_{cond}$ through concatenation. This form $p\left(z_{t}|y_{cond}\right)$ can be implemented with a conditional denoising autoencoder $\epsilon_{\theta}\left(z_{t},t,y_{cond}\right)$. A Unet-style structure [37], whose components include an encoder and a decoder, severs as the denoising network. Based on feature-conditioning pairs, we then learn the conditional LDM via
@@
The definition of a symbols can be refer to the equation (3).
IV-C Detection Head
After generating the feature map $F^{\prime}_{i}$, we decode them into the detection layer $\Phi_{Dec}\left(\cdot\right)$ to obtain final object detections $\mathcal{O}_{i}$. The classification output is the confidence score of being an vehicle or background for each anchor box, the regression output is $\left(x,y,z,w,l,h,\theta\right)$, representing the centroid coordinates, size and yaw of the anchor boxes, respectively.
IV-D Training details and loss function
Following common practice [35, 38, 39], we separate training into two distinct stages. First, we train a perception compression model to provide a low-dimensional representation space. Then, we train the latent diffusion model in this space. In the detection task, we use the cross entropy and weighted smooth L1 loss for the classification $\mathcal{L}_{cls}$ and regression $\mathcal{L}_{reg}$. The total loss is the weighted sum of the diffusion loss in Equation (6) and the detection loss.
@@ IV-D Training details and loss function
V Experiment Result
We validate our CoDiff on both simulated and real-world scenarios. The task of the experiments on the three datasets is point-cloud-based 3D cooperative object detection. Following the literature, the detection performance are evaluated by using Average Precision (AP) at Intersection-over-Union (IoU) thresholds of 0.50 and 0.70.
V-A Datasets
DAIR-V2X [24]. DAIR-V2X is a large-scale vehicle-infrastructure cooperative perception dataset containing over 100 scenes and 18,000 data pairs, featuring two agents: vehicle and road-side unit (RSU), capturing simultaneous data from infrastructure and vehicle sensors at an equipped intersection as an autonomous vehicle passes through.
V2XSet [22]. V2XSet is built with the co-simulation of OpenCDA [42] and CARLA [43]. It is a large-scale simulated dataset designed for Vehicle-to-Infrastructure (V2X) communication. The dataset consists of a total of 11,447 frames, and the train/validation/test splits are 6,694/1,920/2,833, respectively.
OPV2V [23]. OPV2V is a large-scale dataset specifically designed for vehicle-to-vehicle (V2V) communication. It is jointly developed using the CARLA and OpenCDA simulation tools[43]. It includes 12K frames of 3D point clouds and RGB images with 230K annotated 3D boxes.
V-B Implementation Details
We keep PointPillar [1] as LiDAR encoder backbone. It converts the LiDAR point cloud into voxels with the resolution to 0.4m for both height and width. We train an autoencoder to compress the features to an 8-dimensional space with a 32x compression rate. The diffusion model is trained using the Adam optimizer with a learning rate of 0.0002 and weight decay of 0.001. The training is conducted over 20 epochs with a total of 500 steps, adding diffusion conditions at each step. Finally, the object detection model is fine-tuned for 10 epochs. To simulate noisy error and time delay, we follow the noisy settings in RoCo [17] and the delay setting in CoBEVFlow [22] during the training process. We add Gaussian noise $N\left(0,\sigma^{2}_{t}\right)$ on $x,y$ and $N\left(0,\sigma^{2}_{r}\right)$ on $\theta$, where $x,y,\theta$ are the coordinates of the 2D centers of a vechicle and the yaw angle of accurate global poses. The time delay is set to 100 ms. In the training process, the conditions are sequentially added to the Unet according to the steps. All models are trained on six NVIDIA RTX 2080Ti GPUs.
V-C Quantitative evaluation
To validate the overall performance of CoDiff in 3D object detection, we compare it with seven state-of-the-art methods on the three datasets. For a fair comparison, all models take the same 3D point clouds as input data. All methods use the same feature encoder based on PointPillars [1]. Table I shows the AP at IoU threshold of 0.5 and 0.7 in DAIR-V2X, V2XSet and OPV2V dataset. We see that CoDiff significantly outperforms the previous methods at various noise levels across the three datasets and the leading gap is larger when the noise level is higher. Fig. 3 shows the detection performances of the proposed CoDiff and the baseline methods under varying levels of time delay on DAIR-V2X. We see that the proposed CoDiff achieves the best performance at all time delay settings, as shown by the red line in the graph.

### v3

@@ IV Our Proposed Method
To improve this situation, we propose CoDiff, a generative model specifically for multi-agent 3D object detection. As illustrated in Fig. 3, the overall architecture of CoDiff comprises four major components: a single agent detection module, a pose calibration module (PCM), a time calibration module (TCM) and a detection head. Fig. 3 shows the overall architecture of CoDiff. It consists of two key ideas. Firstly, PCM constructs local graphs and exploits graph similarity to establish inter-object correspondences, followed by graph optimization for pose refinement. Secondly, TCM employs a conditional diffusion model to filter time delay noise, thereby generating more reliable multi-agent features. The fused features are finally fed into the detection head for high-precision 3D object detection.
Figure 4: Graph construction and similarity.
IV-A Pose Calibration Module
The input point cloud is represented as a tensor of size $\left(n\times 4\right)$, where $n$ indicates the total number of points. Each point contains its intensity value along with its 3D coordinates in the global reference frame. Furthermore, the system estimates the detection uncertainty by computing the variance distributions of both centroid coordinates and heading angles for each bounding box. All agents utilize a unified 3D object detector, denoted as $\Phi_{Enc}$, and every agent can independently and accurately detect objects within its own range.
After completing the initial detection, each agent $j$ transmits a message $\mathcal{M}_{j\rightarrow i}=(F_{j},\mathcal{B}_{j},\xi_{j})$ to the ego-agent $i$, where $F_{j}$ denotes the feature map, $\mathcal{B}_{j}$ represents the set of detected bounding boxes, and $\xi_{j}=\left(x_{j},y_{j},\theta_{j}\right)$ specifies the agent’s pose, where $(x_{j},y_{j})$ is the 2D position and $\theta_{j}$ the yaw angle relative to the heading angle. Upon receiving messages from all other agents, the ego-agent $i$ sequentially processes each message to perform graph-based pose calibration.
The task can be formally defined as follows: given two sets of information $\left\{\mathcal{B}_{i},\xi_{i}\right\}$ and $\left\{\mathcal{B}_{j},\xi_{j}\right\}$, we need to establish correspondences between the two sets of bounding boxes, as shown in Fig. 4(a). To achieve this, we formulate the task as a bipartite graph matching problem [36], where optimal matches are determined through graph similarity measures. The aforementioned procedure can be formally expressed as:
$\mathcal{A}^{*}_{ij}$ $=\{(p,q)\in\mathcal{B}_{i}\times\mathcal{B}_{j}\mid S(p,q)\geq\tau_{1}\}$ (2a)
$q$ $=\mathcal{A}_{ij}(p)$ (2b)
Where $p\in\mathcal{B}_{i}$ and $q\in\mathcal{B}_{j}$ denote detected objects by agents $i$ and $j$ respectively. $\mathcal{A}_{ij}$ defines a bijective mapping between the object sets, associating each $p$ with its corresponding $q=\mathcal{A}_{ij}(p)$. The matching quality is evaluated through a graph similarity measure $S(p,q)$. A threshold $\tau_{1}$ is a predefined similarity threshold for reliable correspondences.
Graph construction. The proposed PCM module is graph-guided. As illustrated in Fig. 4(b), for each object $p\in\mathcal{B}_{i}$, we construct a star-shaped graph $\mathcal{G}_{p}$ (shown in blue) centered at $p$, where each node represents a feature vector composed of the 3-DoF pose from the object’s bounding box. Similarly, for every object $q\in\mathcal{B}_{j}$, we build a corresponding star-shaped graph $\mathcal{G}_{q}$ (shown in red).
Initial Matching. To initiate the matching process, we first transform the bounding boxes $\mathcal{B}_{j}$ from the coordinate system of agent $j$ into the coordinate frame of agent $i$. Then a distance-based strategy is designed to carry out the initial association. Specifically, for a pair of candidate graphs $\mathcal{G}_{p}$ and $\mathcal{G}_{q}$, we compute the spatial distance between their central nodes $p$ and $q$, denoted as $dis(p,q)$. If this distance is below a predefined threshold $\tau_{2}$, the pair $\mathcal{G}_{p}$ and $\mathcal{G}_{q}$ is considered potentially corresponding and will be included in the matching candidates. It is worth noting that multiple objects from $\mathcal{B}_{j}$ may satisfy this distance constraint with respect to node $p$. In such cases, we select the one with the minimum distance as the initial match, formulated as
$\mathcal{A}_{ij}(p)=\underset{p\in\mathcal{B}_{i},q\in\mathcal{B}_{j}}{\mathrm{arg}\min}\ \ dis\left(p,q\right);\ where\ dis\left(p,q\right)\leq\tau_{2}$ (3)
Graph Matching. Logically, when two detections $p$ and $q$ correspond to the same object, their associated subgraphs $\mathcal{G}_{p}$ and $\mathcal{G}_{q}$ should have high similarity. Taking this into account, we design two similarity metrics, edge similarity and distance similarity, to jointly evaluate the proximity between graphs $\mathcal{G}{p}$ and $\mathcal{G}{q}$, thus determining whether $p$ and $q$ originate from the same object.
  • Edge Similarity. In the graph $\mathcal{G}_{p}$, each edge $e_{pm}$, which connects the central node $p$ to its neighboring node $m$, encodes the relative pose transformation $\mathbf{T}_{pm}\in\mathbb{R}^{3\times 3}$. This transformation $\mathbf{T}_{pm}$ is a $3\times 3$ matrix computed based on the detection outputs of objects $p$ and $m$. Analogously, in the graph $\mathcal{G}_{q}$, the edge $e_{qn}$ connecting node $q$ to its neighbor $n$ encodes the relative pose transformation $\mathbf{T}_{qn}$.
    When $p$ and $q$ correspond to the same object (i.e., $\mathcal{G}_{p}$ and $\mathcal{G}_{q}$ exhibit high structural similarity), their edge transformations should satisfy the consistency condition: $\mathbf{T}_{qn}$ should be consistent with $\mathbf{T}_{pm}$. We establish the following criterion to assess the consistency of the edge between $e_{pm}$ and $e_{qn}$:
    $l(e_{pm},e_{qn})=\exp\left(-\|\mathbf{T}_{pm}(\mathbf{T}_{qn})^{-1}-\mathbf{I}\|_{F}\right)$ (4)
    where $\|\cdot\|_{F}$ denotes the Frobenius norm, $\mathbf{I}$} denotes the identity matrix. The overall edge consistency score is computed as:
    $\begin{gathered}S_{\text{edge}}(p,q)=\frac{1}{|N_{p}|}\sum_{m\in N_{p}}l(e_{pm},e_{qn})\\ s.t.\ \ \ n={\mathcal{A}}_{ij}\left(m\right)\end{gathered}$ (7)
    where $N_{p}$ denotes the leaf nodes in $\mathcal{G}_{p}$ matched with those in $\mathcal{G}_{q}$, $n$ is the counterpart of $m$ from ${\mathcal{A}}_{ij}$.
  • Distance Similarity. Beyond edge similarity, the degree of correspondence between two graphs can be further assessed through distance similarity. The rationale is that when the central nodes $p$ and $q$ represent the same object, the spatial distance of their associated detection boxes should remain minimal, even under noisy measurements or in highly congested scenes. Accordingly, we define the distance similarity between $p$ and $q$ as:
    $S_{dis}\left(p,q\right)=\exp\left(-dis\left(p,q\right)\right)$ (8)
    where $dis(.,.)$ computes the Euclidean distance between bounding box centroids, the overall graph similarity is:
    $S(p,q)$ $=$ $S_{edge}(q,p)+\lambda S_{dis}(p,q)$ (9)
    where $\lambda$ is the hyperparameter for balancing.
Finally, the overall similarity $S$ is used to solve the maximization problem in Equation (9) to obtain the optimal matching. This is treated as a bipartite graph problem, with one set containing bounding boxes from agent $i$ and the other from agent $j$, using previously defined similarity $S(p,q)$ as edge weights. The Kuhn-Munkres algorithm [37] efficiently finds the solution, which is applied iteratively across all agents until every pairwise matching is completed.
Pose Calibration. After completing the object matching, we obtain a one-to-one correspondence between the elements of the object detection box set $\mathcal{B}_{i}$ and $\mathcal{B}_{j}$, As illustrated in Figure 5(a). we construct a pose graph $\mathcal{G}=(\mathcal{V}_{\text{agent}},\mathcal{V}_{\text{object}},\mathcal{E})$ for all agents and their detected objects in the scene, using these correspondences to adjust agent poses. In this pose graph, $\mathcal{V}_{\text{agent}}$ represents all agents $N$, $\mathcal{V}_{\text{object}}$ contains all detected objects, and $\mathcal{E}$ encodes detection relationships between agents and objects. Each node is associated with a pose. The pose of the agent $j$ is denoted as $\xi_{j}$, and the pose of the object $k$ is denoted as $\chi_{k}$.
Consider a matched pair $(p,q)$ where $\mathcal{A}_{ij}(p)=q$. Let $\mathbf{T}_{ip}$ denote the relative transformation of object $p$ measured from agent $i$’s coordinate frame, while $\mathbf{T}_{jq}$ represents the relative transformation of object $q$ observed by agent $j$. From agent $i$’s perspective, we establish the relationship:
$\mathbf{T}_{ip}=\mathbf{E}_{i}^{-1}\mathbf{X}_{p}\Rightarrow\mathbf{X}_{p}=\mathbf{E}_{i}\mathbf{T}_{ip}$ (10)
where $\mathbf{E}_{i}$ is the transformation matrix constructed from agent $i$’s pose $\xi_{i}$. Similarly, we derive $\mathbf{X}_{q}=\mathbf{E}_{j}\mathbf{T}_{jq}$ for agent $j$’s observations. The optimization variables in this framework are the agent and object poses $\{\xi_{j},\chi_{k}\}$. The objective is to achieve pose consistency where matched objects converge to identical coordinates:
$\chi_{p}=\chi_{q}=\chi_{k}$ (11)
Figure 5(b) demonstrates our residual error formulation approach. We define the error functions as:
$\mathbf{e}_{ik}$ $=(\mathbf{E}_{i}\mathbf{T}_{ip})^{-1}\mathbf{X}_{k}=\mathbf{T}_{ip}^{-1}\mathbf{E}_{i}^{-1}\mathbf{X}_{k}$ (12)
$\mathbf{e}_{jk}$ $=(\mathbf{E}_{j}\mathbf{T}_{jq})^{-1}\mathbf{X}_{k}=\mathbf{T}_{jq}^{-1}\mathbf{E}_{j}^{-1}\mathbf{X}_{k}$ (13)
$\mathbf{e}_{ji}$ $=\mathbf{T}_{ji}^{-1}\mathbf{E}_{j}^{-1}\mathbf{E}_{i}$ (14)
The overall optimization objective is formulated as:
$\left\{\bm{\xi}^{\prime}_{j},\bm{\chi}^{\prime}_{k}\right\}=\underset{\left\{\bm{\xi}_{j},\bm{\chi}_{k}\right\}}{\arg}\min\sum_{\left(j,k\right)}{\left(\bm{e}_{ik}^{T}\mathbf{\Omega}_{ik}\bm{e}_{ik}\!\!+\!\!\bm{e}_{jk}^{T}\mathbf{\Omega}_{jk}\bm{e}_{ik}\!\!+\!\!\bm{e}_{ij}^{T}\mathbf{\Omega}_{ij}\bm{e}_{ij}\right)}$ (15)
The information matrix $\mathbf{\Omega}=\mathrm{diag}(\sigma_{x}^{-2},\sigma_{y}^{-2},\sigma_{\theta}^{-2})\in\mathbb{R}^{3\times 3}$ encodes the detection uncertainty associated with agent $j$’s bounding box estimates, which are included in the detection outputs. Applying the Levenberg-Marquardt algorithm [38] to solve (15), we obtain the optimized poses $\{\bm{\xi}_{j}^{\prime}\}_{j=1}^{N}$ for agents and $\{\bm{\chi}_{k}^{\prime}\}_{k=1}^{K}$ for objects. The process converges when the global consistency error is minimized.
Finally, the corrected relative pose from agent $j$ to agent $i$ is represented as $\xi^{\prime}_{j\rightarrow i}=\xi^{-1}_{i}\circ\xi^{\prime}_{j}$. Based on this corrected relative pose $\xi^{\prime}_{j\rightarrow i}$, the message $\mathcal{M}_{j\rightarrow i}$ can be aligned accordingly, so that the feature map $\tilde{F}_{j}$ from agent $j$ is transformed into the ego coordinate frame, making it consistent with agent $i$’s own feature $\mathbf{F}_{i}$.
Figure 5: Pose calibration and the residue errors.
IV-B Time Calibration Module
Upon completing pose calibration, it is necessary to integrate the features from multiple agents to derive the final result. Existing fusion strategies [24, 18] rely on autoregressive or attention-based Transformer models. Nevertheless, we observe that in collaborative detection, these methods overlook the fact that time delay is also one of the causes of feature misalignment, thereby leading to fused features with harmful noise. In TCM, We employed diffusion models for feature fusion, replacing the existing feature fusion methods based on regression models and attention mechanisms. We treat the time delays between multiple agents as a unified noise to be learned and handled.
@@ IV-B Time Calibration Module
Despite their denoising advantage, diffusion models require costly function evaluations during high-resolution feature sampling, leading to significant computational overhead. To address this, we perform sampling in a lower-dimensional latent space, where diffusion models can operate within a perceptually equivalent but computationally efficient representation. This strategy, previously validated in high-resolution image synthesis [35], motivates the design of our perception compression module. This autoencoder comprises an encoder and a decoder, focusing on compressing and reconstructing input tensors along the channel dimension. The encoder employs a series of convolutional layers, batch normalization, and ReLU activation functions [39] to compress high-dimensional input tensors into a lower-dimensional latent space. The decoder, using a similar structure, aims to reconstruct a low-dimensional latent representation back to the original dimensionality. This autoencoder effectively reduces computational overhead while preserving key features of the data.
@@
Precisely, given a feature map $F_{i}\in\mathbb{R}^{H\times W\times C}$ obtained from PCM, where $H,W,C$ represents height, width and channels, respectively. The encoder $\mathcal{E}$} encodes $F_{i}$ into a latent representation $z={\mathcal{E}}\left(F_{i}\right)$ using a compression rate $\tau$, and the decoder $\mathcal{D}$ reconstructs the feature from the latent space using the same compression rate. Therefore, the process of compressing the features $F_{i}$ into the latent space is:
@@
With the perceptual compression model, the diffusion model can now be trained in this lower-dimensional, computationally more efficient latent space. Our goal is to generate noise-free features $F^{\prime}_{i}$ (Step (1c)) through this latent space and the decoder $\mathcal{D}$, which should represent the aggregation of features from multiple agents. We can decode the latent representation of features $F^{\prime}_{i}$ from the latent space back to the original feature space with a single pass through $\mathcal{D}$ during training.
Conditional Denoising Diffusion Module. Diffusion models are generative models inspired by nonequilibrium thermodynamics [41, 42], designed to learn a data distribution $P\left(x\right)$ by gradually denoising a normally distributed variable. The diffusion forward process transforms an initial data sample $x_{0}$ into a noisy sample $x_{t}$ at time $t$ by introducing noise that is determined by the noise variance schedule $\beta_{t}$. For convenience, we denote a series of constants: $\alpha_{t}:=1-\beta_{t},\bar{\alpha}_{t}:=\prod_{s=1}^{t}a_{s}$. The forward process to generate a noisy sample $x_{t}$ from $x_{0}$ can be defined by
where $\mathcal{N}$ denotes a Gaussian distribution and $\mathbf{I}$ is an identity matrix. The neural network is then trained to reverse this diffusion process. This network can be interpreted as an equally weighted sequence of denoising autoencoders $\epsilon_{\theta}\left(x_{t},t\right);t=1,...,T$, which are trained to predict a denoised variant of their input $x_{t}$, where $x_{t}$ is a noisy version of the input $x$. The corresponding objective can be simplified to
Diffusion models can learn the diverse data distribution in multiple domains. We aim to leverage this advantage and design more suitable generative conditions to address the issues of time delays between agents in collaborative perception.
@@
The Design of Condition. Inspired by the literature [35], diffusion models can leverage conditional mechanisms $p\left(x|y\right)$ to control the target generation process, thus improving the accuracy of the model. We find that in multi-agent collaborative perception, the information from individual agents can be naturally used as conditional input to generate the final fused features. Thus, the generation of multi-agent features can be controlled by manipulating the input conditions $y$.
@@
In practice, we select the information on the characteristics of the individual agents N1N-1 as condition $y_{cond}$. These informations are obtained by generating features $F_{j}$ through point cloud encoders, which are then transmitted to the Ego agent $i$, denoted as $\mathcal{M}_{j\rightarrow i}$. We strictly follow the standard DM model to add noise, while the difference is that we employ condition-modulated denoising, which is shown in Figure 3. By progressively denoising the samples, we expect the conditional diffusion model to serve as a more flexible conditional generator, effectively suppressing time-delay-induced noise and thereby facilitating the learning of fine-grained object representations, such as precise boundaries and highly detailed shapes, particularly under noisy conditions.
@@
Given the noisy feature $z_{t}$ and the condition $y_{cond}=\{\mathcal{M}_{j\rightarrow i}\}$ at time step $t$, $z_{t}$ is further encoded and interacts with $y_{cond}$ through concatenation. This form $p\left(z_{t}|y_{cond}\right)$ can be implemented with a conditional denoising autoencoder $\epsilon_{\theta}\left(z_{t},t,y_{cond}\right)$. A Unet-style structure [43], whose components include an encoder and a decoder, severs as the denoising network. Based on feature-conditioning pairs, we then learn the conditional LDM via
@@
The definition of a symbols can be referred to the equation (19). After generating the feature map $F^{\prime}_{i}$, we decode them into the detection layer $\Phi_{Dec}\left(\cdot\right)$ to obtain the final object detections $\mathcal{O}_{i}$.
IV-C Training details and loss function
Following common practice [35, 44], the training process is divided into two stages. First, we train a perception compression model to construct a low-dimensional representation space. Second, a conditional diffusion model is trained within this space. For optimization, we adopt cross-entropy loss for classification ($\mathcal{L}_{cls}$) and weighted smooth L1 loss for regression ($\mathcal{L}_{reg}$). The total loss is the weighted sum of the diffusion loss in Equation (20) and the detection loss.
@@ IV-C Training details and loss function
V Experiment Result
We validate our CoDiff on both simulated and real-world scenarios. Following the literature, the detection performance are evaluated by using Average Precision (AP) at Intersection-over-Union (IoU) thresholds of 0.50 and 0.70.
V-A Datasets
DAIR-V2X [26]. DAIR-V2X is a large-scale vehicle-infrastructure cooperative perception dataset containing over 100 scenes and 18,000 data pairs, featuring two agents: vehicle and road-side unit (RSU), capturing simultaneous data from infrastructure and vehicle sensors at an equipped intersection as an autonomous vehicle passes through. V2XSet [24]. V2XSet is built with the co-simulation of OpenCDA [47] and CARLA [48]. It is a large-scale simulated dataset designed for Vehicle-to-Infrastructure (V2X) communication. The dataset consists of a total of 11,447 frames, and the train/validation/test splits are 6,694/1,920/2,833, respectively. OPV2V [25]. OPV2V is a large-scale dataset specifically designed for vehicle-to-vehicle (V2V) communication. It is jointly developed using the CARLA and OpenCDA simulation tools[48]. It includes 12K frames of 3D point clouds and RGB images with 230K annotated 3D boxes.
V-B Implementation Details
We keep PointPillar [1] as LiDAR encoder backbone. Converts the LiDAR point cloud into voxels with a resolution of 0.4 m for both height and width. In bipartite graph matching, we set the similarity threshold $\tau_{1}=0.5$ for Eq.(2) and $\lambda=1$ for Eq.(9). In pose graph optimization, the Levenberg-Marquardt algorithm[38] is used to solve the least squares optimization problem, with the maximum number of iterations set to 1000. We train an auto-encoder to compress the features to an 8-dimensional space with a 32x compression rate. The diffusion model is trained using the Adam optimizer with a learning rate of 0.0002 and weight decay of 0.001. The training is conducted over 20 epochs with a total of 500 steps, adding diffusion conditions at each step. Finally, the object detection model is fine-tuned for 10 epochs. To simulate pose inconsistencies and time delays, we follow the noisy settings in RoCo [19] and the delay setting in CoBEVFlow [24] during the training process. We add Gaussian noise $N\left(0,\sigma^{2}_{t}\right)$ on $x,y$ and $N\left(0,\sigma^{2}_{r}\right)$ on $\theta$, where $x,y,\theta$ are the coordinates of the 2D centers of a vechicle and the yaw angle of accurate global poses. The time delay is set to 100 ms. All models are trained on six NVIDIA RTX 2080Ti GPUs.
V-C Quantitative evaluation
To validate the overall performance of CoDiff in 3D object detection, we compare it with seven state-of-the-art methods on the three datasets: F-Cooper [45], V2VNet [46], Self-ATT [25], V2X-ViT [24], CoBEVFlow [21] , CoAlign [23] and RoCo [19]. For a fair comparison, all models take the same 3D point clouds as input data. All methods use the same feature encoder based on PointPillars [1]. For a fair comparison, all models take the same 3D point clouds as input data. All methods use the same feature encoder based on PointPillars [1]. Table I shows the AP at IoU threshold of 0.5 and 0.7 in DAIR-V2X, V2XSet and OPV2V dataset. We see that CoDiff significantly outperforms the previous methods at various noise levels across the three datasets and the leading gap is larger when the noise level is higher.
We further investigate the impact of the time delay in the range [0, 400] ms. Fig. 7 shows the detection performances of the proposed CoDiff and the baseline methods under varying levels of time delay on DAIR-V2X. We see that the proposed CoDiff achieves the best performance at all time delay settings, as shown by the red line in the graph. In the case of a time delay of 200ms, our approach achieves an improvement of 2. 6% (82. 0% $vs.$ 79. 4%) and 0. 5% (59. 5% $vs.$ 59. 0%) over RoCo for AP@0.5/0.7. As the delay increases, all methods experience varying degrees of performance degradation, however, even under a 400 ms delay, our approach still achieves superior detection accuracy.